# Is Your Paper Being Reviewed by an LLM? A New Benchmark Dataset and Approach for Detecting AI Text in Peer Review

### v1

@@ A.1.3 Entropy
Similar to the Rank score, the Entropy score for a text is determined by averaging the entropy values of each word, conditioned on its preceding context  Gehrmann et al. (2019).
@@ A.1.4 GLTR
The Entropy score, like the Rank score, is computed by averaging the entropy values of each word within a text, considering the preceding context  Gehrmann et al. (2019).
@@ A.1.5 MFD
The Multi-level Fine-grained Detection (MFD)  Wu and Xiang (2023) framework enhances text detection by combining statistical, semantic, and linguistic features at the sentence level. It first extracts low-level statistical features like readability and author style to quantify sentence structure. Simultaneously, high-level semantic differences are captured using an encoder with contrastive learning to distinguish LLM-generated text from human-written content. Additionally, advanced LLMs analyze the full text, extracting deep linguistic features related to lexicon, grammar, and syntax for more precise detection.
@@ A.1.6 Loglikelihood
This method utilizes a language model to compute the token-wise log probability. Specifically, given a text, the log probability of each token is averaged to produce a final score. A higher score indicates a greater likelihood that the text is machine-generated  Solaiman et al. (2019).
@@ A.1.7 LogRank
Unlike the Rank metric, which relies on absolute rank values, the Log-Rank score is derived by applying a logarithmic function to the rank value of each word  Mitchell et al. (2023b).
@@ A.1.8 Rank
The Rank score is calculated by determining the absolute rank of each word in a text based on its preceding context. The final score is obtained by averaging the rank values across the text. A lower score suggests a higher probability that the text was machine-generated  Gehrmann et al. (2019).
@@ A.1.9 DetectLLM-LLR
This approach integrates Log-Likelihood and Log-Rank scores, leveraging their complementary properties to analyze a given text  Su et al. .
@@ A.1.10 FastDetectGPT
This method assesses changes in a model’s log probability function when small perturbations are introduced to a text. The underlying idea is that LLM-generated text often resides in a local optimum of the model’s probability function. Consequently, minor perturbations to machine-generated text typically result in lower log probabilities, whereas perturbations to human-written text may lead to either an increase or decrease in log probability  Mitchell et al. (2023b).
@@ A.2.1 NTNU-D
It is a fine-tuned classification model based on the RoBERTa-base model, and three sizes of the bloomz-models  Sivesind and Winje (2023)
@@ A.2.2 ChatGPT-D
The ChatGPT Detector  Guo et al. (2023) is designed to differentiate between human-written text and content generated by ChatGPT. It is based on a RoBERTa model that has been fine-tuned for this specific task. The authors propose two training approaches: one that trains the model solely on generated responses and another that incorporates both question-answer pairs for joint training. In our evaluation, we adopt the first approach to maintain consistency with other detection methods.
@@ A.2.3 OpenAI-D and RADAR-D
The OpenAI Detector  Solaiman et al. (2019) are models fine-tuned on RoBERTa to identify outputs generated by GPT-2. Specifically, it was trained using text generated by the largest GPT-2 model (1.5B parameters) and is capable of determining whether a given text is machine-generated.
@@ A.2.4 MAGE-D
MAGE (MAchine-GEnerated text detection)  Li et al. (2024) is a large-scale benchmark designed for detecting AI-generated text. It compiles human-written content from seven diverse writing tasks, including story generation, news writing, and scientific writing. Corresponding machine-generated texts are produced using 27 different LLMs, such as ChatGPT, LLaMA, and Bloom, across three representative prompt types.
@@ Appendix B Dataset Details
B.1 Dataset File Structure
The calibration, test, and extended sets are in separate directories. Each directory contains subdirectories for different models that were used to generate AI peer review samples. In each model’s subdirectory, you will find multiple CSV files, with each file representing peer review samples of a specific conference. The directory and file structure are outlined below.
@@ B.1 Dataset File Structure
B.2 CSV File Content
CSV files may differ in their column structures across conferences and years. These differences are due to updates in the required review fields over time as well as variations between conferences. See Table 5 for review fields of individual conferences.
@@ B.2 CSV File Content
Table 5: Required fields in the review templates for each conference.
B.3 Dataset Sample Numbers per Conference Year
@@
Table 6: Entire set sample size, including both human and AI reviews. They are exactly balanced.
@@
Table 7: Test set sample size, including both human and AI reviews. They are exactly balanced.
@@
Table 8: Calibration set sample size, including both human and AI reviews. They are exactly balanced.
@@ Appendix C Additional results
C.1 Calibration using ICLR + NeurIPS reviews
Our main results in Table 3 of Section 5.1 utilized ICLR review from our calibration set to calibrate each detection method. This simulates the scenario in which some of the reviews in the test set are "out-of-domain" in the sense that they belong to a different conference than the reviews used for calibration. In Table 9, we provide additional results for the same evaluation setting as before, but using both ICLR and NeurIPS reviews for calibration (i.e., fully “in domain”). We generally see similar trends regarding relative performance between methods as before, with the exception that the Binoculars method achieves slightly higher detection rates than our Anchor method for Gemini reviews. This suggests that existing methods such as Binoculars may be more sensitive to the use of in-domain data during calibration.
@@ C.1 Calibration using ICLR + NeurIPS reviews
Table 9: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR and NeurIPS reviews from our calibration set at different target FPRs. Best TPRs are in bold.
C.2 Additional results for test set: Llama and Qwen Detection Results
Table 10 is organized similarly to Table 3, but it presents results for Llama and Qwen reviews. Both tables use the same set of thresholds for each method, which were calibrated using ICLR reviews generated by GPT-4o, Gemini, and Claude along with their matching human-written reviews.
@@ C.2 Additional results for test set: Llama and Qwen Detection Results
Table 10: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR reviews from our calibration set at different target FPRs. Best TPRs are in bold.
C.3 Experimental Results on Full Dataset
We test existing AI text generation text detection models on our entire dataset (i.e., the test set + the extended set). The results are shown in Tables 11 for ICLR reviews and 12 for NeurIPS reviews.
@@ C.3 Experimental Results on Full Dataset
Table 11: Actual FPR and TPR calculated from the ICLR reviews at varying detection thresholds, which are calibrated using the ICLR calibration dataset at different target FPRs.
@@
GPT BinocularsMetric 0.3% 26.1% 0.6% 45.3% 0.9% 54.1%
GLTRMetric 0.3% 1.3% 5.1% 31.9% 10.6% 52.6%
Llama BinocularsMetric 0.3% 98.9% 0.6% 99.3% 0.9% 99.4%
GLTRMetric 0.3% 76.3% 0.9% 91.0% 1.4% 93.9%
Table 12: Actual FPR and TPR calculated from the NeurIPS reviews at varying detection thresholds, which are calibrated using the ICLR calibration dataset at different target FPRs.
@@
Table 13: Examples of differences identified in human analysis of human and AI-written peer reviews
Appendix D Additional analyses
D.1 Examples from human analysis of differences between human and AI-written peer reviews
Table 13 provides examples of the issues identified in our qualitative analysis of human and AI-written peer reviews. In general, we observe that GPT-4o reviews lack references to specific details in the paper, lack references to specific prior work, and contain overly generic criticisms. See Section 6.1 for additional discussion.
D.2 Comparison of numeric scores assigned by human and AI reviewers
While Section 6.2 focused on the misalignment between human and AI peer reviews from three commercial LLMs (GPT-4o, Gemini, and Claude) from the NeurIPS2022 review samples, this section presents the corresponding results for two open-source LLMs (Llama and Qwen), as shown in Figure 3. The main findings from GPT-4o, Gemini, and Claude also hold for these two open-source models, with one notable difference: Llama and Qwen exhibit an even larger divergence in Presentation scores than Claude, the most overly-positive one amongst the commercial LLMs across all categories. In terms of Contribution scores, the evaluations from Llama and Qwen were similar to those of Claude.

### v2

@@ A.1.3 Entropy
Similar to the Rank score, the Entropy score for a text is determined by averaging the entropy values of each word, conditioned on its preceding context  [31].
@@ A.1.4 GLTR
The Entropy score, like the Rank score, is computed by averaging the entropy values of each word within a text, considering the preceding context  [31].
@@ A.1.5 MFD
The Multi-Feature Detection (MFD) method [57] detects AI-generated text using four features: log-likelihood, log-rank, entropy, and LLM deviation.
@@ A.1.6 Loglikelihood
This method utilizes a language model to compute the token-wise log probability. Specifically, given a text, the log probability of each token is averaged to produce a final score. A higher score indicates a greater likelihood that the text is machine-generated  [27].
@@ A.1.7 LogRank
Unlike the Rank metric, which relies on absolute rank values, the Log-Rank score is derived by applying a logarithmic function to the rank value of each word  [58].
@@ A.1.8 Rank
The Rank score is calculated by determining the absolute rank of each word in a text based on its preceding context. The final score is obtained by averaging the rank values across the text. A lower score suggests a higher probability that the text was machine-generated  [31].
@@ A.1.9 DetectLLM-LLR
This approach integrates Log-Likelihood and Log-Rank scores, leveraging their complementary properties to analyze a given text  [36].
@@ A.1.10 FastDetectGPT
This method assesses changes in a model’s log probability function when small perturbations are introduced to a text. The underlying idea is that LLM-generated text often resides in a local optimum of the model’s probability function. Consequently, minor perturbations to machine-generated text typically result in lower log probabilities, whereas perturbations to human-written text may lead to either an increase or decrease in log probability  [58].
@@ A.2.1 NTNU-D
It is a fine-tuned classification model based on the RoBERTa-base model, and three sizes of the bloomz-models  [59]
@@ A.2.2 ChatGPT-D
The ChatGPT Detector  [20] is designed to differentiate between human-written text and content generated by ChatGPT. It is based on a RoBERTa model that has been fine-tuned for this specific task. The authors propose two training approaches: one that trains the model solely on generated responses and another that incorporates both question-answer pairs for joint training. In our evaluation, we adopt the first approach to maintain consistency with other detection methods.
@@ A.2.3 OpenAI-D and RADAR-D
OpenAI Detector [27] is a fine-tuned RoBERTa model designed to identify outputs generated by GPT-2. Specifically, it was trained using text generated by the largest GPT-2 model (1.5B parameters) and is capable of determining whether a given text is machine-generated.
@@ A.2.4 MAGE-D
MAGE (MAchine-GEnerated text detection)  [60] is a large-scale benchmark designed for detecting AI-generated text. It compiles human-written content from seven diverse writing tasks, including story generation, news writing, and scientific writing. Corresponding machine-generated texts are produced using 27 different LLMs, such as ChatGPT, LLaMA, and Bloom, across three representative prompt types.
@@ Appendix B Dataset Details
B.1 Dataset
Our dataset is hosted on Hugging Face and publicly available at https://huggingface.co/datasets/IntelLabs/AI-Peer-Review-Detection-Benchmark. It is released under the Intel OBL Internal R&D Use License Agreement (https://huggingface.co/datasets/IntelLabs/AI-Peer-Review-Detection-Benchmark/blob/main/LICENSE.md).
B.2 Dataset File Structure
The calibration, test, and extended sets are in separate directories. Each directory contains subdirectories for different models that were used to generate AI peer review samples. In each model’s subdirectory, you will find multiple CSV files, with each file representing peer review samples of a specific conference. Each file follows the naming convention: "<conference>.<subset>.<LLM>.csv". The directory and file structure are outlined below.
@@ B.2 Dataset File Structure
B.3 CSV File Content
CSV files may differ in their column structures across conferences and years. These differences are due to updates in the required review fields over time as well as variations between conferences. See Table S1 for review fields of individual conferences.
Table S1: Required fields in the review templates for each conference.
@@ B.3 CSV File Content
B.4 Dataset Sample Numbers per Conference Year
@@
Table S2: Entire set sample size, including both human and AI reviews. They are exactly balanced.
@@
Table S3: Test set sample size, including both human and AI reviews. They are exactly balanced.
@@
Table S4: Calibration set sample size, including both human and AI reviews. They are exactly balanced.
@@ Appendix C Additional results
C.1 Uncertainty Estimates via Bootstrap Resampling
To estimate the uncertainty of our main experimental results (TPR and FPR reported in Table 2), we perform bootstrap resampling with replacement using $N=100$ resamples. For each method and evaluation setting, we compute the standard deviation (SD) of the TPR and FPR across the 100 bootstrap replicates. While 100 resamples provide only a coarse estimate of variability, this level of resampling was chosen to balance computational cost with the need to quantify uncertainty. As shown in Table S5, the standard deviations are small relative to the large performance gaps between methods, indicating the robustness of our main findings.
Table S5: Actual true positive rate (TPR) and false positive rate (FPR), $\pm$ their standard deviation, computed over 100 bootstrap resamples corresponding to the main results in Table 2.
Target FPR: 0.1% 0.5% 1%
FPR TPR FPR TPR FPR TPR
GPT-4o Reviews Anchor 0.1 $\pm$ 0.02 63.5 $\pm$ 0.34 0.5 $\pm$ 0.03 83.7 $\pm$ 0.33 1.0 $\pm$ 0.04 88.8 $\pm$ 0.31
Binoculars 0.2 $\pm$ 0.03 17.1 $\pm$ 0.25 0.6 $\pm$ 0.05 33.6 $\pm$ 0.31 1.0 $\pm$ 0.06 45.2 $\pm$ 0.34
MAGE-D 0.1 $\pm$ 0.03   2.3 $\pm$ 0.11 0.6 $\pm$ 0.05   8.8 $\pm$ 0.21 1.3 $\pm$ 0.07 14.7 $\pm$ 0.24
s5 0.1 $\pm$ 0.03   0.1 $\pm$ 0.02 0.9 $\pm$ 0.07   7.2 $\pm$ 0.20 1.7 $\pm$ 0.09 17.5 $\pm$ 0.29
MFD 0.2 $\pm$ 0.03   0.1 $\pm$ 0.02 0.8 $\pm$ 0.07   6.0 $\pm$ 0.15 1.6 $\pm$ 0.09 15.6 $\pm$ 0.24
GLTR 0.1 $\pm$ 0.02   0.1 $\pm$ 0.02 0.4 $\pm$ 0.05   1.9 $\pm$ 0.08 1.1 $\pm$ 0.08   5.7 $\pm$ 0.14
DetectGPT 0.1 $\pm$ 0.02   0.1 $\pm$ 0.02 0.6 $\pm$ 0.05   1.1 $\pm$ 0.07 1.2 $\pm$ 0.08   2.3 $\pm$ 0.10
Gemini Reviews Anchor 0.2 $\pm$ 0.01 59.7 $\pm$ 0.32 0.8 $\pm$ 0.03 80.3$\pm$ 0.33 1.3 $\pm$ 0.04 86.5 $\pm$ 0.31
Binoculars 0.2 $\pm$ 0.03 61.5 $\pm$ 0.32 0.6 $\pm$ 0.05 78.0 $\pm$ 0.28 1.0 $\pm$ 0.07 85.5 $\pm$ 0.23
s5 0.0 $\pm$ 0.01   0.2 $\pm$ 0.03 0.5 $\pm$ 0.04   9.6 $\pm$ 0.21 1.1 $\pm$ 0.08 19.4 $\pm$ 0.27
MFD 0.1 $\pm$ 0.02   0.5 $\pm$ 0.04 0.4 $\pm$ 0.05   8.9 $\pm$ 0.17 1.1 $\pm$ 0.07 18.8 $\pm$ 0.25
FastDetectGPT 0.1 $\pm$ 0.02   1.1 $\pm$ 0.08 0.5 $\pm$ 0.04   5.8 $\pm$ 0.18 1.1 $\pm$ 0.06 10.3 $\pm$ 0.21
GLTR 0.2 $\pm$ 0.03   0.5 $\pm$ 0.04 0.8 $\pm$ 0.06   5.2 $\pm$ 0.13 1.8 $\pm$ 0.09 12.4 $\pm$ 0.23
DetectGPT 0.1 $\pm$ 0.02   0.4 $\pm$ 0.05 0.5 $\pm$ 0.05   3.5 $\pm$ 0.12 1.2 $\pm$ 0.08   7.0 $\pm$ 0.18
MAGE-D 0.1 $\pm$ 0.03   0.4 $\pm$ 0.04 0.6 $\pm$ 0.06   3.3 $\pm$ 0.12 1.3 $\pm$ 0.08   7.0 $\pm$ 0.17
Loglikelihood 0.0 $\pm$ 0.01   0.0 $\pm$ 0.00 0.3 $\pm$ 0.04   0.1 $\pm$ 0.02 0.5 $\pm$ 0.05   1.0 $\pm$ 0.07
Claude Reviews Anchor 0.1 $\pm$ 0.02 59.6 $\pm$ 0.34 0.5 $\pm$ 0.04 75.8 $\pm$ 0.34 1.0 $\pm$ 0.05 81.8 $\pm$ 0.32
Binoculars 0.2 $\pm$ 0.04 43.5 $\pm$ 0.33 0.6 $\pm$ 0.05 65.8 $\pm$ 0.33 1.0 $\pm$ 0.07 77.0 $\pm$ 0.29
s5 0.0 $\pm$ 0.01   0.1 $\pm$ 0.02 0.2 $\pm$ 0.03   7.6 $\pm$ 0.18 0.5 $\pm$ 0.05 17.5 $\pm$ 0.23
MFD 0.0 $\pm$ 0.01   0.2 $\pm$ 0.03 0.1 $\pm$ 0.03   6.8 $\pm$ 0.18 0.4 $\pm$ 0.05 16.5 $\pm$ 0.24
DetectGPT 0.1 $\pm$ 0.02   0.5 $\pm$ 0.05 0.6 $\pm$ 0.05   5.3 $\pm$ 0.16 1.2 $\pm$ 0.06 11.1 $\pm$ 0.22
GLTR 0.0 $\pm$ 0.01   0.0 $\pm$ 0.01 0.2 $\pm$ 0.03   0.5 $\pm$ 0.05 0.6 $\pm$ 0.05   1.8 $\pm$ 0.10
C.2 Calibration using ICLR + NeurIPS reviews
Our main results in Table 2 of Section 4.1 utilized ICLR review from our calibration set to calibrate each detection method. This simulates the scenario in which some of the reviews in the test set are "out-of-domain" in the sense that they belong to a different conference than the reviews used for calibration. In Table S6, we provide additional results for the same evaluation setting as before, but using both ICLR and NeurIPS reviews for calibration (i.e., fully “in domain”). We generally see similar trends regarding relative performance between methods as before, with the exception that the Binoculars method achieves slightly higher detection rates than our Anchor method for Gemini reviews. This suggests that existing methods such as Binoculars may be more sensitive to the use of in-domain data during calibration.
Table S6: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR and NeurIPS reviews from our calibration set at different target FPRs. Best TPRs are in bold.
@@ C.2 Calibration using ICLR + NeurIPS reviews
C.3 Additional Llama and Qwen Detection Results
Table S7 is organized similarly to Table 2, but presents results for Llama and Qwen reviews. Both tables use the same set of thresholds for each method, which were calibrated using ICLR reviews generated by GPT-4o, Gemini, and Claude along with their matching human-written reviews.
Reviews generated by open-source LLMs (Table S7) show somewhat different trends compared to reviews generated by commercial LLMs (Table 2). For both Llama and Qwen reviews, the Binoculars method achieves near-perfect detection performance. While the Anchor method ranks second for Qwen reviews, it is not as performant on Llama reviews. This is surprising, as Llama reviews appear easier to detect. For instance, at a target FPR of 0.1%, more than 6.5 times as many methods achieve a TPR above 10% compared to GPT-4o, Gemini, and Claude reviews, and about twice as many compared to Qwen reviews. One possibility is that the semantic similarity between the three anchor reviews generated by higher-quality LLMs (GPT-4o, Gemini, and Claude) and Llama reviews is low enough to overlap with that between the anchor reviews and human reviews, potentially blurring the decision boundary. Although the underlying causes of these differences warrant further investigation, these findings are less central to our study since most LLM users are likely to rely on commercial models due to their ease of use and superior capabilities.
Table S7: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR reviews from our calibration set at different target FPRs. Best TPRs are in bold. Detection methods are ordered by their TPR at a target FPR of 0.1%, and those that failed to achieve 10% TPR at a target FPR of 1% are omitted.
@@ C.3 Additional Llama and Qwen Detection Results
C.4 Experimental Results for Full Dataset
We test existing AI text generation text detection models on our entire dataset (i.e., the test set + the extended set). Results are provided in Tables S8 and S9 for ICLR and NeurIPS reviews (rspeectively).
Table S8: Actual FPR and TPR calculated from the ICLR reviews at varying detection thresholds, which are calibrated using the ICLR calibration dataset at different target FPRs.
@@ C.4 Experimental Results for Full Dataset
Table S9: Actual FPR and TPR calculated from the NeurIPS reviews at varying detection thresholds, which are calibrated using the ICLR calibration dataset at different target FPRs.
@@
GPT Binoculars 0.3% 26.1% 0.6% 45.3% 0.9% 54.1%
GLTR 0.3% 1.3% 5.1% 31.9% 10.6% 52.6%
Llama Binoculars 0.3% 98.9% 0.6% 99.3% 0.9% 99.4%
GLTR 0.3% 76.3% 0.9% 91.0% 1.4% 93.9%
C.5 Computation Time
We report the computation time of each detection method. Each method was evaluated on 100 samples using a single NVIDIA RTX A6000 GPU, repeated 20 times to compute the mean and standard deviation. The Anchor method is an exception: it does not use a GPU and relies on sequential API calls to the OpenAI service, which could potentially be optimized for faster execution (e.g., parallelizing API calls). Figure S1 summarizes the results.
Figure S1: Computation time (in seconds) for processing 100 samples. Each method was repeated 20 times to compute the mean and standard deviation. All methods were run on a single NVIDIA RTX A6000 GPU, except for Anchor, which used sequential API calls without GPU acceleration.
Appendix D Additional analyses
D.1 Voting Mechanism for the Anchor Embedding Method
Intuitively, the anchor approach performs best when the anchor embeddings are generated using the same model that produced the test review (source LLM). However, in real-world scenarios, the source LLM is typically unknown (a situation commonly referred to as “black-box” detection scenario). To address this challenge, we propose a voting-based technique. Specifically, we generate multiple anchor embeddings using different types of LLMs (anchor LLMs). For each anchor embedding, we compute the Score (Eq.4) and derive the corresponding label assignment (Eq.5). If at least one anchor embedding assigns a positive label, the final label is positive. Otherwise, the final label is negative. In our experiments, we used three anchor reviews for voting—each generated by GPT-4o, Gemini, and Claude, respectively.
D.2 Assessing Dataset Robustness to Prompt Variation (Supplementary)
This appendix provides supporting materials for the prompt sensitivity analysis described in Section 5.2 of the main text. The results presented here include quantitative evaluations of detection performance under prompt variation (Table S10) and a visualization of review embeddings to assess stylistic consistency across prompt types (Figure S2). Together, these results support the conclusion that our prompting strategy yields stable and generalizable outputs across different prompt formulations.
Table S10: Actual FPR and TPR on the withheld ICLR2021 test set. The upper section shows results for reviews generated using the score-aligned prompt (used in our main dataset), and the lower section shows results for reviews generated using the alternative, archetype-based prompt. In both cases, the same set of thresholds, calibrated on the calibration set, is applied. Detection performance remains largely consistent across prompt styles, indicating that prompt variation does not substantially degrade model effectiveness.
Target FPR: 0.1% 0.5% 1%
FPR TPR FPR TPR FPR TPR
Score-aligned Anchor 0.2 86.9 0.8 96.4 1.6 98.1
Binoculars 0.1 7.6 0.3 21.9 0.6 32.6
s5 0.1 0.0 0.7 3.8 1.6 12.5
MFD 0.1 0.0 0.8 2.8 1.5 10.8
MAGE-D 0.0 0.0 0.1 1.1 0.3 2.4
GLTR 0.1 0.0 0.4 0.3 1.1 1.8
Archetype Anchor 0.2 83.4 0.8 94.8 1.6 96.8
Binoculars 0.1 7.8 0.3 18.1 0.6 27.5
s5 0.1 0.0 0.8 6.2 1.6 14.9
MFD 0.1 0.0 0.8 4.5 1.7 12.8
MAGE-D 0.0 0.1 0.1 0.6 0.3 1.7
GLTR 0.1 0.0 0.3 0.0 0.8 0.8
Figure S2: t-SNE visualization of sentence embeddings from AI-generated reviews in the ICLR2021 test set. Blue points represent reviews generated using the main score-aligned prompt, and orange points represent those from the alternative archetype-based prompt. The substantial overlap between the two distributions suggests that prompt variation does not cause major shifts in model outputs. Embeddings were computed using OpenAI’s text-embedding-3-small model; t-SNE was performed with 2 output dimensions and a perplexity of 30.
D.3 Examples from human analysis of differences between human and AI-written peer reviews
Table S11 provides examples of the issues identified in our qualitative analysis of human and AI-written peer reviews. In general, we observe that GPT-4o reviews lack references to specific details in the paper, lack references to specific prior work, and contain overly generic criticisms. See Section 5.4 for additional discussion.
Table S11: Examples of differences identified in human analysis of human and AI-written peer reviews
@@ D.3 Examples from human analysis of differences between human and AI-written peer reviews
D.4 Comparison of numeric scores assigned by human and AI reviewers
While Section 5.5 focused on the misalignment between human and AI peer reviews from three commercial LLMs (GPT-4o, Gemini, and Claude) from the NeurIPS2022 review samples, this section presents the corresponding results for two open-source LLMs (Llama and Qwen), as shown in Figure S3. The main findings from GPT-4o, Gemini, and Claude also hold for these two open-source models, with one notable difference: Llama and Qwen exhibit an even larger divergence in Presentation scores than Claude, the most overly-positive one amongst the commercial LLMs across all categories. In terms of Contribution scores, the evaluations from Llama and Qwen were similar to those of Claude.